OpenStudy (anonymous):

EXAM TOMORROW PLEASE HELP Minimize objective function p(x,y)=2x+3y subject to these constraints: 2x+y 21 x /> 0 y /> 0 I used as greater than or equal to, etc.

OpenStudy (anonymous):

first draw st. lines 2x+y=14 and x+3y=21

satellite73 (satellite73):

\[x\geq 0,y\geq 0\]means you are in quadrant 1

satellite73 (satellite73):

we also need \(2x+y\leq 14\) and \(x+3y\geq 21\] graphed, and the points where they intersect

satellite73 (satellite73):

oops \[x+3y\geq 21\]

satellite73 (satellite73):

do you know how to ind the intersection?